OpenStudy (anonymous):

Choose the correct first step(s) to solve the system using substitution. y = 5x and 2x + 3y = 34

OpenStudy (anonymous):

assuming simultaneous, called y = 5x Eq.1 and 2x + 3y = 34 Eq.2 Multiply Eq.1 by 3 to get 3y = 15x, then substitute into Eq.2 Which yields 17x = 35 and therefore x = 35/17 Then substitute x = 35/17 back into original Eq.1 to give y = 175/17

OpenStudy (anonymous):

2x + 3(5x) = 34 2x + 15x = 34 17x = 34 x = 34/17

OpenStudy (anonymous):

Thank you guys so much : )